OpenStudy (anonymous):

Using the Ambiguous Law of Sines, solve the following triangle. You will have two solutions. In Triangle ABC: a=9 c=15 angleA=35˚ Please solve step by step?

OpenStudy (anonymous):

"Ambiguous"? law of sines; a/sinA = b/sinB = c/sinC where a is the side opposite angle A etc.

OpenStudy (solomonzelman):

2 solutions? Well the law of sines states, \[\huge\color{red}{ \frac{Sin(A)}{a}=\frac{Sin(B)}{b} =\frac{Sin(C)}{c} }\] Perhaps the hyp doesn't have to be integer, and then side c can be either hyp or one of the legs. Only this way there are TWO solutions.

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

i dont understand how there are 2

OpenStudy (solomonzelman):

SO you would have to draw 2 shapes in this case. See?

OpenStudy (anonymous):

ohhh i understand

OpenStudy (solomonzelman):

Yeah, the first one is 3-4-5 pythagorean triple dilated by the scale factor of 3. and the second would give you the hyp which will be irrational. I think you have to solve each of those shapes fully to get each answer.

OpenStudy (anonymous):

i can't get the answer

OpenStudy (solomonzelman):

That's what we have the Sin law for. \[\huge\color{blue}{ \frac{Sin~A}{a} =\frac{Sin~B}{b} =\frac{Sin~C}{c} }\] a sin of angle divided by the side it forms, EQUALS a sin of another angle divided by the side it forms. (IN SAME TRIANGLE)

OpenStudy (solomonzelman):

For example, in the first shape, you want to find angle B. you would use a^2+b^2=c^2 to find b, and then plug in b into \[\frac{Sin~B}{b~~~~(which~~~you~~~found)} =\frac{Sin~35}{9} \] plug in b and solve for angle B. Can you solve for angle B now?

OpenStudy (anonymous):

yes i can